For the first six years, the loan value of a Corvette is determined by the formula
V = – 4000a + 50000, where V is the value and a is the age of the car in years. What was its original value? What is its value when it is 5 years old?

Answer by Abbey(339)

The original value is determined when the age of the car is zero.
V=-4000(0)+50000
V=50,000
When it is 5 years old
V=-4000(5)+50000
V=-20000+50000=30000
so the value of the car is 30,000 after 5 years